do with a ti 84 and explain how you got your answers

the reaction time X (in minutes) of a certain chemical process follows a uniform probability distribution with 5 < X < 10.

draw a graph of the density curve.

what is the probability that the reaction time is between 6 and 8 minutes?

what is the probability that the reaction time is between 5 and 8 minutes?

what is the probability that the reaction time is less than 6 minutes?
Answer #1

(these are simple calculation for which you require simple calculaton,(Also uniform distribution is not available on ti-84)

Distribution Plot Uniform, Lower=5, Upper=10 0.20 Density 0.05 0.00 5

probability that the reaction time is between 6 and 8 minutes =P(6 <X<8) =(8-6)/(10-5)=0.4

probability that the reaction time is between 5 and 8 minutes =P(5 <X<8) =(8-5)/(10-5)=0.6

probability that the reaction time is less than 6 minutes =P(X<6) =(6-5)/(10-5)=0.2